OpenStudy (anonymous):

can someone please help!!!! will medal ...An expression is shown below: f(x) = –16x2 + 60x + 16 Part A: What are the x-intercepts of the graph of the f(x)? Show your work. Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work. Part C: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph.

OpenStudy (anonymous):

I dont want someone to just tell me the answer i want to learn how to do this type of problem

OpenStudy (anonymous):

@jim_thompson5910 could you help when you help someone you explain it very well :)

OpenStudy (freckles):

Part A: To find the x-intercepts, set y=f(x) to 0 and solve the equation for x. You could quadratic formula or factoring if it factorable or completing the square.

OpenStudy (anonymous):

ok so it would be 0=-16x^2 +60x+16

OpenStudy (freckles):

yep

OpenStudy (freckles):

you could make those numbers easier on the eyes if you divide both side by 4 giving you 0=-4x^2+15x+4

OpenStudy (freckles):

this expression on the right hand side of the equation is factorable if you want to factor it

OpenStudy (anonymous):

ok i have that write next in my notes and then im lost

OpenStudy (freckles):

ok well do you know the quadratic formula?

OpenStudy (freckles):

\[0=ax^2+bx+c \implies x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

no i dont know how to do that yet :(

OpenStudy (freckles):

ok I guess we will try factoring if you have ax^2+bx+c you can look for two factors that have product ac and sum b. that means multiply to be ac and add up to b

OpenStudy (freckles):

you have -4x^2+15x+4 a=-4 b=15 c=4 ac=-16 So we want to find two numbers that multiply to be -16 and add up to be 15

OpenStudy (freckles):

can you think of two such numbers

OpenStudy (freckles):

pairs of numbers that multiply to be -16: 1(-16) -1(16) 2(-8) -2(8) 4(-4) do any of these pairs of factors add up to be 15?

OpenStudy (anonymous):

16 and -1

OpenStudy (freckles):

perfect

OpenStudy (freckles):

so 15=16-1

OpenStudy (freckles):

so we can write our middle term 15x as 16x-1x why does this help?

OpenStudy (freckles):

we have unlock the perfect sum such that we can factor by grouping

OpenStudy (freckles):

\[-4x^2+15x+4 \\ \\-4x^2+16x-1x+4 \\ (-4x^2+16x)+ -1x+4 \\ (-4x^2+16x)+(-1x+4)\]

OpenStudy (freckles):

now looking at the first group there what is the gcf of the terms inside the first ( )

OpenStudy (anonymous):

4

OpenStudy (freckles):

anything else be sure to look at the number of factors of x in each term

OpenStudy (anonymous):

sorry 4x

OpenStudy (freckles):

cool stuff

OpenStudy (freckles):

so we will factor 4x from the first group

OpenStudy (freckles):

\[-4x^2+15x+4 \\ \\-4x^2+16x-1x+4 \\ (-4x^2+16x)+ -1x+4 \\ (-4x^2+16x)+(-1x+4) \\ 4x(-1x+4)+(-1x+4)\]

OpenStudy (anonymous):

so 4x(-x+4)

OpenStudy (freckles):

you should now see two terms 4x(-1x+4) and (-1x+4) and in these terms you should see the common factor (-1x+4)

OpenStudy (anonymous):

wow your fast

OpenStudy (freckles):

\[-4x^2+15x+4 \\ \\-4x^2+16x-1x+4 \\ (-4x^2+16x)+ -1x+4 \\ (-4x^2+16x)+(-1x+4) \\ 4x(-1x+4)+(-1x+4) \\ 4x(-1x+4)+1(-1x+4)\] I'm going to put a 1 next to our next group because people forget the 1 is there sometimes

OpenStudy (freckles):

now you are going to factor out the (-1x+4) and yes you can it (-x+4) if you want :)

OpenStudy (freckles):

Rember ba+ca=(b+c)a or a(b+c) example (kitty)(dog)+(kitty)(fish)=(kitty)(dog+fish)

OpenStudy (freckles):

so can show me what you think the factored expression looks like

OpenStudy (anonymous):

(4x+1) (-x+4) (-x+4)

OpenStudy (freckles):

you have one too many factors

OpenStudy (anonymous):

-x+4

OpenStudy (anonymous):

need to be taken out

OpenStudy (freckles):

\[4x(-x+4)+1(-x+4) \\ \color{blue}{4x}\color{red}{(-x+4)} \color{blue}{+1 }\color{red}{(-x+4)} \\ \color{red}{(-x+4)}\color{blue}{(4x+1)}\] right

OpenStudy (freckles):

so we did have an equation earlier we were only playing with the right hand expression to get it in factored form

OpenStudy (freckles):

so we had \[0=-16x^2+60x+16 \\ 0=-4x^2+15x+4 \\ 0=(-x+4)(4x+1)\]

OpenStudy (freckles):

now you set both factors equal to 0 because if you have something times something = 0 then at least one of those somethings is zero

OpenStudy (freckles):

so solve the equations: -x+4=0 , 4x+1=0

OpenStudy (anonymous):

so the first on i have x=4 or -x=-4 the second one i got x= -1/4

OpenStudy (freckles):

(x,y) is an ordered pair remember we let y be 0 <-- this was to find the x-intercepts So we know our answer should in the form (x,0)

OpenStudy (freckles):

So your answers for the first part is: The x-intercepts are (4,0) and (-1/4,0)

OpenStudy (freckles):

Did you have any questions on the first part?

OpenStudy (anonymous):

no you explained where i was having trouble

OpenStudy (freckles):

If you don't know the formula for the vertex... Then you must be able to complete the square to find the vertex... \[y=-16x^2+60x+16 \] I'm going to assume we don't know the formula and just find the vertex form for this parabola which entails the completing the square thing I was talking about \[y=-16x^2+60x+16 \\ y=(-16x^2+60x)+16\] We are going to look at these first two terms only for now... We are going to factor the number tagged onto the x^2 from both terms in the ( )

OpenStudy (freckles):

that is we are going to factor out -16 from both terms inside the ( )

OpenStudy (freckles):

you know -16/-16=1 \[y=(-16x^2+60x)+16 \\ y=\frac{-16}{-16}(-16x^2+60x)+16 \\ \\ y=-16(\frac{1}{-16})(-16x^2+60x)+16 \\ y=-16(\frac{-16x^2+60x}{-16})+16 \\ y=-16(\frac{-16}{-16}x^2+\frac{60}{-16}x)+16 \] I wrote it like this on purpose because a lot of students have trouble seeing how to factor numbers out of other numbers where those numbers don't divide the other number evenly that is for example in this question you do not see a factor of -16 in 60

OpenStudy (freckles):

anyways let's go ahead and reduce those fractions inside the ( )

OpenStudy (freckles):

\[y=-16(1x^2+\frac{-15}{4}x)+16 \\ y=-16(x^2+\frac{-15}{4}x)+16\] let me know if you understand everything up til this point

OpenStudy (anonymous):

ok

OpenStudy (freckles):

time to add something inside the grouping symbol so we can complete the square now now whatever we add into that grouping symbol we are going to have remember that thing is being multiplied by the -16 on the outside so whatever we put in we had to take out I will show you what I mean in a bit

OpenStudy (freckles):

to complete the square you need to have the coefficient of x^2 is 1 we have that now we just did that... second we take the number in front of x.... divide by 2 and then square and that is how we figure out what number to add inside the ( )

OpenStudy (freckles):

example \[x^2+kx+(\frac{k}{2})^2=(x+\frac{k}{2})^2\]

OpenStudy (freckles):

\[x^2+\frac{-15}{4}x+.... =(x+......)^2\]

OpenStudy (freckles):

so we take the number in front of x: divide it by 2...(or you can say multiply by 1/2) what do you get when you do this?

OpenStudy (anonymous):

-15/8 i think

OpenStudy (freckles):

you think right

OpenStudy (anonymous):

yeah

OpenStudy (freckles):

ok then square... \[x^2+\frac{-15}{4}x+(\frac{-15}{8})^2 =(x+\frac{-15}{8})^2 \\ \text{ just like in the above formula I wrote } \\ x^2+kx+(\frac{k}{2})^2=(x+\frac{k}{2})^2 \text{ note: where } k=\frac{-15}{4}\]

OpenStudy (freckles):

ok let's go back to our problem... \[y=-16(1x^2+\frac{-15}{4}x)+16 \\ y=-16(x^2+\frac{-15}{4}x)+16 \\ y=-16(x^2+\frac{-15}{4}x+?)+16+16?\] that ? we are fixing to plug in is (-15/8)^2 but before we do that I just want you to see that we are actually adding in 0... notice that first ? is being multiplied by -16 so we are actually putting in -16? so we have to put in 16? notice-16?+16? is =0

OpenStudy (freckles):

\[y=-16(x^2+\frac{-15}{4}x+(\frac{-15}{8})^2) +16+16(\frac{-15}{8})^2\]

OpenStudy (freckles):

now again the whole reason we did this so we could group the terms with x in it together and then write it as a square

OpenStudy (freckles):

\[y=-16(x+\frac{-15}{8})^2+16+16(\frac{-15}{8})^2\]

OpenStudy (freckles):

you can add the like terms there that is you can simplify: \[16+16(\frac{-15}{8})^2\]

OpenStudy (freckles):

once you do that it should be really clear this is in vertex form \[y=a(x-h)^2+k \text{ where vertex is } (h,k)\] also if a is negative then (h,k) is a max if a is positive then (h,k) is a min

OpenStudy (freckles):

I will give you time to digest before saying anything else

OpenStudy (anonymous):

so then would you find out what (-15/8)^2 is

OpenStudy (anonymous):

which is 225/64

OpenStudy (freckles):

yes \[16+16(\frac{225}{64})=...\]

OpenStudy (anonymous):

would i multiply 16 times 225 and 16 time 64 next

OpenStudy (freckles):

nope

OpenStudy (freckles):

a(b/c) is equal to ab/c not ab/(ac)

OpenStudy (freckles):

\[16+\frac{16 \cdot 225}{64}\]

OpenStudy (anonymous):

16+3600/64

OpenStudy (freckles):

and the end you should get 289/4 after adding those fractions and reducing

OpenStudy (freckles):

or you can reduce that one fraction then add

OpenStudy (freckles):

\[y=-16(x-\frac{15}{8})^2+\frac{289}{4}\]

OpenStudy (freckles):

so you should be able to identify the vertex

OpenStudy (freckles):

remember y=a(x-h)^2+k tells us the vertex is (h,k) this also tells us if a is positive then (h,k) is min and that if a is negative then (h,k) is max

OpenStudy (freckles):

please let me know what you think the vertex is

OpenStudy (anonymous):

so because 16 is negative it is max

OpenStudy (freckles):

since -16=a this means a is negative so the vertex will be a max

OpenStudy (freckles):

so yes I would change what you said to "so because -16 is negative it is max"

OpenStudy (anonymous):

ok :}

OpenStudy (anonymous):

Thank you for taking your time to help me on this problem im going to have to finish it up tomorrow but thank you so much for your help you were great !!!!!!!!

OpenStudy (freckles):

The cool thing about this problem is that all of the hard math is out of the way: You are only left with the following things comparing y=-16(x-15/8)^2+289/4 to y=a(x-h)^2+k to identify (h,k) which is the vertex And then graphing by plotting your x-intercepts and vertex and connecting the dots forming a parabola

OpenStudy (freckles):

Anyways goodnight and good luck